# The Uniform Mordell–Lang Conjecture

### v3

@@ 3.2. Family version
Take $b\in\iota_{\mathbf{H}^{\circ}}(S)(\overline{\mathbb{Q}})$. Use the notation from (3.3) with $A=A_{b}$. For any $\underline{P}\in(\pi^{\mathrm{univ},[m]})^{-1}(b)(\overline{\mathbb{Q}})=A_{b}^{m}(\overline{\mathbb{Q}})$, we have
(3.8) $(\iota^{[m]}|_{\mathscr{X}^{[m]}_{S}})^{-1}(\underline{P})=(\iota_{A_{b}}^{[m]}|_{\mathscr{X}_{r,d}(A_{b})^{[m]}\times_{\mathbf{H}_{r,d}(A_{b})}V})^{-1}(\underline{P}),$
where $V=(\iota_{\mathbf{H}^{\circ}}|_{S})^{-1}(b)\subseteq\iota_{\mathbf{H}^{\circ}}^{-1}(b)=\mathbf{H}_{r,d}(A_{b})^{\circ}$. Thus there exists $\underline{P}$ such that $(\iota^{[m]}|_{\mathscr{X}^{[m]}_{S}})^{-1}(\underline{P})$ has dimension 00 by Lemma 3.2.
Thus it remains to prove that ${\mathscr{X}^{[m]}_{S}}$ is irreducible. To do this, by [Gro67, Prop.9.7.8] it suffices to prove that the geometric generic fiber $\mathscr{X}_{\bar{\eta}}$ is irreducible. This is true because $\mathscr{X}_{s}$ is irreducible for all $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ (by definition of $\mathbf{H}^{\circ}$) and [Gro67, Prop.9.7.8].∎
@@ 3.3. Non-degeneracy
We keep the notation from the previous subsection.
Later on, we wish to apply the height inequality of Dimitrov–Gao–Habegger [DGH21, Thm.1.6 and Thm.B.1] and Kühne’s equidistribution result [Küh21, Thm.1] to appropriate abelian schemes and non-degenerate subvarieties to achieve the desired uniform bound. Thus it is fundamental to construct some non-degenerate subvarieties to apply these tools.
For our purpose and inspired by the second-named author’s [Ge21, Prop.3.4], we prove the following proposition. In fact, it is precisely [Ge21, Prop.3.4] adapted to our context of restricted Hilbert schemes, and the main idea of the proof is to apply the first-named author’s [Gao20, Thm.10.1] to the situation considered in Lemma 3.3.
@@
Let $S\subseteq\mathbf{H}^{\circ}$ be an irreducible subvariety. Consider $\mathscr{X}_{S}\subseteq\mathcal{A}_{S}\rightarrow S$.
Use $\bar{\eta}$ to denote the geometric generic point of $S$. Assume that the following hypotheses hold true on the geometric generic fiber $\mathcal{A}_{\bar{\eta}}$ of $\mathcal{A}_{S}/S$:
@@
Then there exists $m_{0}\geq 1$ such that $\mathscr{X}^{[m]}_{S}$ is a non-degenerate subvariety of $\mathcal{A}_{S}^{[m]}$ for each $m\geq m_{0}$.
Non-degenerate subvarieties defined over $\mathbb{C}$ are defined as in [Gao21, Defn.6.1] (we need both descriptions later on). A subvariety defined over $\overline{\mathbb{Q}}$ is said to be non-degeneate if its base change to $\mathbb{C}$ is non-degenerate.
Proof.
We wish to invoke [Gao20, Thm.10.1.(i)] with $t=0$. In fact, as the conventions of [Gao20] is somewhat different from standard terminologies, we will apply the formulation [Gao21, Thm.6.5.(i)].
All hypotheses of [Gao21, Thm.6.5] are satisfied: indeed, hypothesis (a) clearly holds true because we fixed $r\geq 1$ at the beginning of this section, hypothesis (b) holds true by (ii), and hypothesis (c) holds true by (iii). Thus we can apply [Gao21, Thm.6.5.(i)] to the abelian scheme $\mathcal{A}_{S}\rightarrow S$ and the subvariety $\mathscr{X}_{S}$. Therefore $\mathscr{X}^{[m]}_{S}$ is non-degenerate if $m\geq\dim S$ and $\iota^{[m]}|_{\mathscr{X}^{[m]}_{S}}$ is generically finite. By Lemma 3.3 there exists $m_{0}\geq 1$ such that $\iota^{[m]}|_{\mathscr{X}^{[m]}_{S}}$ is generically finite for each $m\geq m_{0}$. So the conclusion holds true with $m_{0}$ replaced by $\max\{m_{0},\dim S\}$. ∎
The following lemma is useful.
@@
Let $S\subseteq\mathbf{H}^{\circ}$ be a (not necessarily irreducible) subvariety. Define the subset
@@
For (i): assume $\mathscr{X}_{\bar{\eta}}$ has $n$ irreducible components. Then $\mathscr{X}_{s^{\prime}}$ has $n$ irreducible components for $s^{\prime}\in U(\overline{\mathbb{Q}})$ with $U$ a Zariski open dense subset of $S^{\prime}$; see [Gro67, Prop.9.7.8]. By definition of $\mathbf{H}^{\circ}$, $\mathscr{X}_{s^{\prime}}$ is irreducible for all $s^{\prime}\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$. Hence $n=1$. So $\mathscr{X}_{\bar{\eta}}$ has $1$ irreducible component, and hence (i) is established.
(ii) clearly holds true.
For (iii): let $C$ be the neutral component of $\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}}(\mathscr{X}_{\bar{\eta}})$, then there exists a quasi-finite dominant morphism $\rho\colon S^{\prime\prime}\rightarrow S^{\prime}$ such that $C$ extends to an abelian subscheme ${\mathcal{C}}$ of $\mathcal{A}_{S^{\prime\prime}}=\mathcal{A}_{S^{\prime}}\times_{S^{\prime}}S^{\prime\prime}\rightarrow S^{\prime\prime}$. So ${\mathcal{C}}_{s^{\prime\prime}}\subseteq\mathrm{Stab}_{\mathcal{A}_{s^{\prime\prime}}}(\mathscr{X}_{s^{\prime\prime}})$ for each $s^{\prime\prime}\in S(\overline{\mathbb{Q}})$. Now that $\rho(S^{\prime\prime})$ contains a Zariski open dense subset of $S^{\prime}$, there exists a point $s^{\prime}\in\rho(S^{\prime\prime})(\overline{\mathbb{Q}})\cap S_{\mathrm{gen}}$ as in the previous paragraph. Take $s^{\prime\prime}\in S^{\prime\prime}(\overline{\mathbb{Q}})$ over $s^{\prime}\in S^{\prime}(\overline{\mathbb{Q}})$, then $\mathcal{A}_{s^{\prime\prime}}$ can be identified with $\mathcal{A}_{s^{\prime}}$ under the natural projection $\mathcal{A}_{S^{\prime\prime}}=\mathcal{A}_{S^{\prime}}\times_{S^{\prime}}S^{\prime\prime}\rightarrow\mathcal{A}_{S^{\prime}}$. Notice that $\mathscr{X}_{s^{\prime\prime}}$ is identified with $\mathscr{X}_{s^{\prime}}$ under the same projection. Thus ${\mathcal{C}}_{s^{\prime\prime}}\subseteq\mathrm{Stab}_{\mathcal{A}_{s^{\prime\prime}}}(\mathscr{X}_{s^{\prime\prime}})=\mathrm{Stab}_{\mathcal{A}_{s^{\prime}}}(\mathscr{X}_{s^{\prime}})$. But $\mathrm{Stab}_{\mathcal{A}_{s^{\prime}}}(\mathscr{X}_{s^{\prime}})$ is finite by definition of $S_{\mathrm{gen}}$ (3.9). So ${\mathcal{C}}_{s^{\prime\prime}}$ is the origin of $\mathcal{A}_{s^{\prime\prime}}$. So ${\mathcal{C}}$ is the zero section of $\mathcal{A}_{S^{\prime\prime}}\rightarrow S^{\prime\prime}$, and hence $C={\mathcal{C}}_{\bar{\eta}}$ is the origin of $\mathcal{A}_{\bar{\eta}}$. Thus we have established (iii). ∎
@@ 4. Applying the height inequality
Let $g$, $l$, $r\leq g$ and $d$ be positive integers. There exist constants $c_{1}^{\prime}=c_{1}^{\prime}(g,l,r,d)>0$, $c_{2}^{\prime}=c_{2}^{\prime}(g,l,r,d)>0$ and $c_{3}^{\prime}=c_{3}^{\prime}(g,l,r,d)>0$ satisfying the following property. For
  • and each irreducible subvariety $X$ of $A$, defined over $\overline{\mathbb{Q}}$, with $\dim X=r$ and $\deg_{L}X=d$, such that $X$ generates $A$,
@@
(4.1) $\Sigma:=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes L_{-}}(x)\leq c_{1}^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}-c_{3}^{\prime}\right\},$
where $L_{-}:=[-1]^{*}L$, satisfies the following property: $\Sigma\subseteq X^{\prime}(\overline{\mathbb{Q}})$ for some Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}(X^{\prime})<c_{2}^{\prime}$.
The key to prove this proposition is to put all $(A,L)$ and $X$ into finitely many families. Briefly speaking, up to some reductions it suffices to prove the proposition for $(A,L)$ principally polarized; then one works with the universal abelian variety over the fine moduli space and consider the restricted Hilbert scheme from (3.6). The core of this section is the result in family Proposition 4.2.
4.1. A result in family
Let $\mathfrak{A}_{g}\rightarrow\mathbb{A}_{g}$ be the universal abelian variety over the fine moduli space of principally polarized abelian varieties with level-$4$-structure. There exists a symmetric relatively ample line bundle $\mathfrak{L}_{g}$ on $\mathfrak{A}_{g}/\mathbb{A}_{g}$ satisfying the following property: for each principally polarized abelian variety $(A,L)$ parametrized by $b\in\mathbb{A}_{g}(\overline{\mathbb{Q}})$, we have $c_{1}(\mathfrak{L}_{g}|_{(\mathfrak{A}_{g})_{b}})=2c_{1}(L)$; see [MFK94, Prop.6.10]. In particular if we identify $A=(\mathfrak{A}_{g})_{b}$, then $c_{1}(\mathfrak{L}_{g}|_{A})=c_{1}(L\otimes L_{-})$ for $L_{-}:=[-1]^{*}L$.
@@ 4.1. A result in family
(4.2) ${\mathscr{X}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{\iota}$${\mathfrak{A}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\mathrm{univ}}}$${\mathbf{H}^{\circ}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota_{\mathbf{H}^{\circ}}}$${\mathbb{A}_{g}}$
where $\pi|_{\mathscr{X}_{\mathbf{H}^{\circ}}}\colon\mathscr{X}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$ is the universal family. Every object in the diagram (4.2) is a variety defined over $\overline{\mathbb{Q}}$, and every morphism is defined over $\overline{\mathbb{Q}}$.
Set ${\mathcal{L}}:=\iota^{*}\mathfrak{L}_{g}^{\otimes 4}$. Then ${\mathcal{L}}$ is relatively very ample on $\mathcal{A}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$.
Fix $\mathfrak{M}$ an ample line bundle on a compactification $\overline{\mathbb{A}_{g}}$ of $\mathbb{A}_{g}$ defined over $\overline{\mathbb{Q}}$. The morphism $\iota_{\mathbf{H}^{\circ}}$ extends to a morphism $\overline{\iota_{\mathbf{H}^{\circ}}}\colon\overline{\mathbf{H}^{\circ}}\rightarrow\overline{\mathbb{A}_{g}}$ for some compactification $\overline{\mathbf{H}^{\circ}}$ defined over $\overline{\mathbb{Q}}$ of $\mathbf{H}^{\circ}$, i.e. $\overline{\mathbf{H}^{\circ}}$ is a projective variety defined over $\overline{\mathbb{Q}}$ which contains $\mathbf{H}^{\circ}$ as a Zariski open dense subset. Set ${\mathcal{M}}:=\overline{\iota_{\mathbf{H}^{\circ}}}^{*}\mathfrak{M}$.
@@
There exist constants $c_{1}^{\prime}=c_{1}^{\prime}(r,d,S)>0$, $c_{2}^{\prime}=c_{2}^{\prime}(r,d,S)>0$ and $c_{3}^{\prime}=c_{3}^{\prime}(r,d,S)>0$ such that the following property holds true. For each $s\in S_{\mathrm{gen}}$, the set
@@
satisfies the following property: $\Sigma_{s}\subseteq X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{2}^{\prime}$.
This proposition is the generalization of Dimitrov–Gao–Habegger’s result for the universal curve [DGH21, Prop.7.1]. The current formulation is closer to the one presented in the first-named author’s survey [Gao21, Prop.7.2]. The proof of Proposition 4.2 is in line with these results; the main new ingredients are Proposition 3.4 (to replace [Gao20, Thm.1.2’] for the non-degeneracy of a new family) and Lemma 4.3 (which generalizes [DGH21, Lem.6.3] to treat subvarieties other than curves).
Proof.
It suffices to prove the proposition for $S$ irreducible.
From now on, assume $S$ is irreducible. We prove this proposition by induction on $\dim S$. The proof of the base step $\dim S=0$ is in fact contained in the induction.
Endow each irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$ with the reduced induced subscheme structure.
Lemma 3.5 allows us to invoke Proposition 3.4 for $\mathscr{X}_{S^{\prime}}\subseteq\mathcal{A}_{S^{\prime}}\rightarrow S^{\prime}$. Let $m_{0}(S^{\prime})>0$ be from Proposition 3.4, and let $m=\max_{S^{\prime}}\{m_{0}(S^{\prime})\}$ with $S^{\prime}$ running over all irreducible components of $S_{\mathrm{gen}}$. Then $\mathscr{X}^{[m]}_{S^{\prime}}$ is a non-degenerate subvariety of $\mathcal{A}_{S^{\prime}}^{[m]}$. Notice that $m=m(S)>0$ depends only on $S$.
Let $\mathscr{X}^{[m],*}_{S^{\prime}}$ be the Zariski open subset of $\mathscr{X}^{[m]}_{S^{\prime}}$ as in [Gao21, Thm.7.1]; it is defined over $\overline{\mathbb{Q}}$ since $\mathscr{X}^{[m]}_{S^{\prime}}$ is, and is dense in $\mathscr{X}^{[m]}_{S^{\prime}}$ since $\mathscr{X}^{[m]}_{S^{\prime}}$ is non-degenerate.
@@
Consider $S^{\prime}\setminus\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})$ endowed with the reduced induced subscheme structure; it has dimension $\leq\dim S^{\prime}-1\leq\dim S-1$. Let $S_{1},\ldots,S_{k}$ be the irreducible components of $\bigcup_{S^{\prime}}S^{\prime}\setminus\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})$ with $S^{\prime}$ running over all irreducible components of $\overline{S_{\mathrm{gen}}}$. Then the set $\{S_{1},\ldots,S_{k}\}$ is uniquely determined by $S$ and $m$. But $m=m(S)>0$ depends only on $S$. So the set $\{S_{1},\ldots,S_{k}\}$ is uniquely determined by $S$.
Let $s\in S_{\mathrm{gen}}(\overline{\mathbb{Q}})$. Then either $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$ or $s\in\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})(\overline{\mathbb{Q}})$ for some irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$.
Case (i) $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$.
For each $i\in\{1,\ldots,k\}$, we can apply the induction hypothesis to $S_{i}$ since $\dim S_{i}\leq\dim S-1$. So we obtain constants $c_{i,1}^{\prime}=c_{i,1}^{\prime}(r,d,S_{i})>0$, $c_{i,2}^{\prime}=c_{i,2}^{\prime}(r,d,S_{i})>0$, and $c_{i,3}^{\prime}=c_{i,3}^{\prime}(r,d,S_{i})>0$ such that for each $s\in S_{i}(\overline{\mathbb{Q}})$, the set

### v4

@@ 3.2. Parameterizing subvarieties in families
Choose some $b\in\iota_{\mathbf{H}^{\circ}}(S)(\overline{\mathbb{Q}})$. We use the notation from (3.3) with $A=A_{b}$. For any $\underline{P}\in(\pi^{\mathrm{univ},[m]})^{-1}(b)(\overline{\mathbb{Q}})=A_{b}^{m}(\overline{\mathbb{Q}})$, we have
(3.8) $(\iota^{[m]})^{-1}(\underline{P})\cap\mathscr{X}_{S}^{[m]}=(\iota_{A_{b}}^{[m]}|_{\mathscr{X}_{r,d}(A_{b})^{[m]}\times_{\mathbf{H}_{r,d}(A_{b})}V})^{-1}(\underline{P}),$
where $V=(\iota_{\mathbf{H}^{\circ}}|_{S})^{-1}(b)\subseteq\iota_{\mathbf{H}^{\circ}}^{-1}(b)=\mathbf{H}_{r,d}(A_{b})^{\circ}$. Thus there exists $\underline{P}$ such that $(\iota^{[m]}|_{\mathscr{X}^{[m]}_{S}})^{-1}(\underline{P})$ has dimension 0 by Lemma 3.2.
The asserted generic finiteness follows if $\mathscr{X}^{[m]}_{S}$} is irreducible. To prove this irreducibility, we use that $\mathscr{X}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$ is flat by construction, and so are the morphisms $\mathscr{X}_{\mathbf{H}^{\circ}}^{[m]}\rightarrow\mathbf{H}^{\circ}$ and ${\mathscr{X}^{[m]}_{S}}\rightarrow S$. Openness of flat morphisms [Gro67, Thm. 2.4.6] implies that every irreducible component of $\mathscr{X}^{[m]}_{S}$} dominates $S$, i.e. its projection contains the generic point of $S$. But since $\mathscr{X}_{s}$ is geometrically irreducible for all $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$, there can be only one such irreducible component of $\mathscr{X}^{[m]}_{S}$} by [Gro67, Prop. 9.7.8]. ∎
@@ 3.3. Non-degeneracy
We keep the notation from the previous subsection, in particular those in (3.5). For later applications of both the height inequality of [DGH21, Thm. 1.6 and B.1] and the equidistribution result [Küh21, Thm. 1], we need to work with appropriate non-degenerate subvarieties in certain families of abelian varieties. Here, we give the main proposition providing such non-degenerate subvarieties, inspired by a result of the second-named author [Ge24, Prop. 3.4] and using a theorem of the first-named author [Gao20, Thm. 10.1].
@@
Let $S\subseteq\mathbf{H}^{\circ}$ be an irreducible subvariety and consider $\mathscr{X}_{S}\subseteq\mathcal{A}_{S}\rightarrow S$. Writing $\bar{\eta}$ for the geometric generic point of $S$, assume that the following hypotheses hold true:
@@
Then there exists $m_{0}=m_{0}(S)\geq 1$ such that $\mathscr{X}^{[m]}_{S}$ is a non-degenerate subvariety of $\mathcal{A}_{S}^{[m]}$ for each $m\geq m_{0}$.
Non-degenerate subvarieties over $\mathbb{C}$ are defined as in [DGH21, Defn. 1.5 and B.4]. A subvariety defined over $\overline{\mathbb{Q}}$ is said to be non-degenerate if its base change to $\mathbb{C}$ is non-degenerate.
Proof of Proposition 3.4.
We wish to invoke [Gao20, Thm. 10.1.(i)] with $t=0$. In fact, as the conventions of [Gao20] deviate a litte bit from standard terminology, we apply it in the formulation of [Gao21, Thm. 6.5.(i)].
All hypotheses of [Gao21, Thm. 6.5] are satisfied: indeed, hypothesis (a) clearly holds true because we fixed $r\geq 1$ at the beginning of this section, hypothesis (b) holds true by (ii), and hypothesis (c) holds true by (iii). Thus we can apply [Gao21, Thm. 6.5.(i)] to the abelian scheme $\mathcal{A}_{S}\rightarrow S$ and the subvariety $\mathscr{X}_{S}$. It yields that $\mathscr{X}^{[m]}_{S}$ is non-degenerate if $m\geq\dim S$ and $\iota^{[m]}|_{\mathscr{X}^{[m]}_{S}}$ is generically finite. By Lemma 3.3 there exists $m_{0}\geq 1$ such that $\iota^{[m]}|_{\mathscr{X}^{[m]}_{S}}$ is generically finite for each $m\geq m_{0}$. So the conclusion holds true with $m_{0}$ replaced by $\max\{m_{0},\dim S\}$. ∎
We conclude with a further lemma useful in later sections.
@@
For a (not necessarily irreducible) subvariety $S\subseteq\mathbf{H}^{\circ}$, we define the subset
@@
To verify the condition (i), we can use again [Gro67, Prop. 9.7.8] as in the proof of Lemma 3.3 above. As there, we simply use the fact that $\mathscr{X}_{s^{\prime}}$ is (geometrically) irreducible for all $s^{\prime}\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$.
The condition (ii) clearly holds true, for example by Chevalley’s Theorem on the upper semicontinuity of the fiber dimension [Gro67, Thm. 13.1.3].
To prove (iii), let $C$ be the identity component of $\mathrm{Stab}_{\mathcal{A}_{\bar{\eta}}}(\mathscr{X}_{\bar{\eta}})$. There exists a quasi-finite dominant morphism $\rho\colon S^{\prime\prime}\rightarrow S^{\prime}$ such that $C$ extends to an abelian subscheme $\mathcal{C}$} of $\mathcal{A}_{S^{\prime\prime}}=\mathcal{A}_{S^{\prime}}\times_{S^{\prime}}S^{\prime\prime}\rightarrow S^{\prime\prime}$. Clearly, ${\mathcal{C}}_{s^{\prime\prime}}\subseteq\mathrm{Stab}_{\mathcal{A}_{s^{\prime\prime}}}(\mathscr{X}_{s^{\prime\prime}})$ for each $s^{\prime\prime}\in S(\overline{\mathbb{Q}})$. Since $\rho(S^{\prime\prime})$ contains a Zariski open dense subset of $S^{\prime}$, there exists a point $s^{\prime}\in\rho(S^{\prime\prime})(\overline{\mathbb{Q}})\cap S_{\mathrm{gen}}$ as in the previous paragraph. Choosing a point $s^{\prime\prime}\in S^{\prime\prime}(\overline{\mathbb{Q}})$ over $s^{\prime}\in S^{\prime}(\overline{\mathbb{Q}})$, the fiber $\mathcal{A}_{s^{\prime\prime}}$ is naturally identified with $\mathcal{A}_{s^{\prime}}$, which also induces an identification of $\mathscr{X}_{s^{\prime\prime}}$ with $\mathscr{X}_{s^{\prime}}$. It follows that ${\mathcal{C}}_{s^{\prime\prime}}\subseteq\mathrm{Stab}_{\mathcal{A}_{s^{\prime\prime}}}(\mathscr{X}_{s^{\prime\prime}})=\mathrm{Stab}_{\mathcal{A}_{s^{\prime}}}(\mathscr{X}_{s^{\prime}})$. By the definition of $S_{\mathrm{gen}}$ (3.9), $\mathrm{Stab}_{\mathcal{A}_{s^{\prime}}}(\mathscr{X}_{s^{\prime}})$ is finite and hence ${\mathcal{C}}_{s^{\prime\prime}}$ is the identity element of $\mathcal{A}_{s^{\prime\prime}}$. Thus $\mathcal{C}$} is the zero section of $\mathcal{A}_{S^{\prime\prime}}\rightarrow S^{\prime\prime}$, and $C={\mathcal{C}}_{\bar{\eta}}$ is the origin of $\mathcal{A}_{\bar{\eta}}$, whence (iii). ∎
@@ 4. Applying the height inequality
Let $g$, $l$, $r\leq g$ and $d$ be positive integers. There exist constants $c_{1}^{\prime}=c_{1}^{\prime}(g,l,r,d)>0$, $c_{2}^{\prime}=c_{2}^{\prime}(g,l,r,d)>0$ and $c_{3}^{\prime}=c_{3}^{\prime}(g,l,r,d)>0$ satisfying the following property: For
  • every polarized abelian variety $(A,L)$ over $\overline{\mathbb{Q}}$ of dimension $g$ and degree $\deg_{L}A=l$, and
  • every irreducible subvariety $X$ of $A$ with $\dim X=r$, $\deg_{L}X=d$ and such that $X$ generates $A$,
@@
(4.1) $\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes[-1]^{*}L}(x)\leq c_{1}^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}-c_{3}^{\prime}\right\},$
is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for a proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}(X^{\prime})<c_{2}^{\prime}$.
@@
4.1. A preliminary result for restricted Hilbert schemes
[…]
For each $1\leq r\leq g$ and each $d\geq 1$, we have the restricted Hilbert scheme $\mathbf{H}^{\circ}:=\mathbf{H}_{r,d}(\mathfrak{A}_{g}/\mathbb{A}_{g})^{\circ}$ from (3.6). Recall that it is a variety defined over $\overline{\mathbb{Q}}$ whose closed points parametrizes all pairs $(X,(A,L))$ of a principally polarized abelian variety $(A,L)$ and an irreducible subvariety $X$ of $A$ with $\dim X=r$ and $\deg_{\mathfrak{L}_{g}^{\otimes 4}|_{A}}X=d$.
@@ 4.1. A preliminary result for restricted Hilbert schemes
[…]
where $\pi|_{\mathscr{X}_{\mathbf{H}^{\circ}}}\colon\mathscr{X}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$ is the universal family. Every object in the diagram (4.2) is a variety over $\overline{\mathbb{Q}}$, and every morphism is defined over $\overline{\mathbb{Q}}$.
Set ${\mathcal{L}}:=\iota^{*}\mathfrak{L}_{g}^{\otimes 4}$. Notice that $\mathcal{L}$} is relatively very ample with respect to $\mathcal{A}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$.
Fix an ample line bundle $\mathfrak{M}$ on a compactification $\overline{\mathbb{A}_{g}}$ of $\mathbb{A}_{g}$. The morphism $\iota_{\mathbf{H}^{\circ}}$ extends to a morphism $\overline{\iota_{\mathbf{H}^{\circ}}}\colon\overline{\mathbf{H}^{\circ}}\rightarrow\overline{\mathbb{A}_{g}}$ for some compactification $\overline{\mathbf{H}^{\circ}}$ of $\mathbf{H}^{\circ}$. We can do these constructions such that $\overline{\mathbf{H}^{\circ}}$ is a projective variety defined over $\overline{\mathbb{Q}}$ which contains $\mathbf{H}^{\circ}$ as a Zariski open dense subset. Set ${\mathcal{M}}:=\overline{\iota_{\mathbf{H}^{\circ}}}^{*}\mathfrak{M}$.
@@
There exist constants $c_{1}^{\prime}=c_{1}^{\prime}(r,d,S)>0$, $c_{2}^{\prime}=c_{2}^{\prime}(r,d,S)>0$ and $c_{3}^{\prime}=c_{3}^{\prime}(r,d,S)>0$ with the following property. For each $s\in S_{\mathrm{gen}}$, the set
@@
is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for a proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{2}^{\prime}$.
This proposition is a generalization of the corresponding result for the universal curve [DGH21, Prop. 7.1], though our formulation here is closer to [Gao21, Prop. 7.2]. Notice that the case $S=\mathbf{H}^{\circ}$ contains the full strength of the proposition. Furthermore, it is the only case we use below in the proof of Proposition 4.4. However, the above formulation is more convenient for the proof by induction.
Proof of Proposition 4.2.
It suffices to prove the proposition for $S$ irreducible, which we assume from now on.
For the proof, we use an induction on $\dim S$. The base case $\dim S=0$ is straightforward as we may choose $c_{3}^{\prime}$ large enough to render $\Sigma_{s}$ empty.
Endow each irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$ with the reduced induced subscheme structure. Lemma 3.5 allows us to invoke Proposition 3.4 for $\mathscr{X}_{S^{\prime}}\subseteq\mathcal{A}_{S^{\prime}}\rightarrow S^{\prime}$. Let $m_{0}(S^{\prime})>0$ be as in Proposition 3.4, and let $m=\max_{S^{\prime}}\{m_{0}(S^{\prime})\}$ with $S^{\prime}$ running over all irreducible components of $S_{\mathrm{gen}}$. Then $\mathscr{X}^{[m]}_{S^{\prime}}$ is a non-degenerate subvariety of $\mathcal{A}_{S^{\prime}}^{[m]}$. Note that $m=m(S)>0$ depends only on $S$.
Let $\mathscr{X}^{[m],*}_{S^{\prime}}$ be $\mathscr{X}^{[m]}_{S^{\prime}}$ deprived of its 0-th degeneracy locus; a definition of the 0-th degeneracy can be found in [Gao20, Defn. 1.6]. By [Gao20, Thm. 1.8], it is a Zariski open subset of $\mathscr{X}^{[m]}_{S^{\prime}}$. It is defined over $\overline{\mathbb{Q}}$ since $\mathscr{X}^{[m]}_{S^{\prime}}$ is and it is dense in $\mathscr{X}^{[m]}_{S^{\prime}}$ because of the non-degeneracy of $\mathscr{X}^{[m]}_{S^{\prime}}$.
@@
Consider $S^{\prime}\setminus\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})$ endowed with the reduced induced subscheme structure; it has dimension $\leq\dim S^{\prime}-1\leq\dim S-1$. Let $S_{1},\ldots,S_{k}$ be the irreducible components of $\bigcup_{S^{\prime}}S^{\prime}\setminus\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})$ with $S^{\prime}$ running over all irreducible components of $\overline{S_{\mathrm{gen}}}$. The set $\{S_{1},\ldots,S_{k}\}$ is uniquely determined by $S$ and $m$. As $m=m(S)>0$ depends only on $S$, it is actually already determined by $S$.
Let $s\in S_{\mathrm{gen}}(\overline{\mathbb{Q}})$. Then either $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$ or $s\in\pi^{[m]}(\mathscr{X}_{S^{\prime}}^{[m],*})(\overline{\mathbb{Q}})$ for some irreducible component $S^{\prime}$ of $\overline{S_{\mathrm{gen}}}$. For the remaining proof of the proposition, we distinguish between these two cases.
Case (i): $s\in\bigcup_{i=1}^{k}S_{i}(\overline{\mathbb{Q}})$.
For each $i\in\{1,\ldots,k\}$, we can apply the induction hypothesis to $S_{i}$ as $\dim S_{i}\leq\dim S-1$. In this way, we obtain constants $c_{i,1}^{\prime}=c_{i,1}^{\prime}(r,d,S_{i})>0$, $c_{i,2}^{\prime}=c_{i,2}^{\prime}(r,d,S_{i})>0$, and $c_{i,3}^{\prime}=c_{i,3}^{\prime}(r,d,S_{i})>0$ such that for each $s\in S_{i}(\overline{\mathbb{Q}})$, the set